x-1=y+1
x+1=2(y-1)

Answer by richwmiller(17219)

subtract one from two after simplifying second
x+1=2y-2
x-1=y-1
0+2=y-1
3=y
x-1=y-1
x-1=2
x=3
check
3+1=2(3-1)
4=2(2)
4=4
ok